OpenStudy (anonymous):

Part 1: Create your own factorable trinomial. Part 2: Explain, in complete sentences, how the trinomial is factored. Part 3: Explain, in complete sentences, the process used to check the factors for accuracy.

OpenStudy (cwrw238):

x^2 - 8x + 12 = 0 this factors to product of 3 binomials of the form ( x +- a(x+-b) to find a and b: we need two numbers whose product is + 12 and whose sum is -8 a little thought gives you -6 and -2 so factors are (x - 6)(x - 2) to check this expand the brackets: x*x -2*x -6*x -6*-2 = x^2 - 2x - 6x + 12 = x^2 -8x + 12 which is our original trinomial

OpenStudy (cwrw238):

* product of 2 binomials